3 blocks are placed in a smooth horizontal surface. the blocks are named A,BandC. A external force F acts on first body. What is the normail reaction on block C?  

Grade:10

1 Answers

pooja nayak
8 years ago
F-R=MAa
R-N=MBa
N=MCa
  =MCF/(MA+MB+MC)
 WHERE R STANDS FOR THE NORMAL FORCE ON B 
AND N STANDS FOR THE NORNAL FORCE ON BLOCK C
a IS THE ACCELERATION OF THE WHOLE SYSTEM